tammy drove 693 miles in 11hrs at the same rate how long would it take her to dive 315 miles?
Found 2 solutions by checkley79, Alan3354:
Answer by checkley79(3341)

683/11=315/X CROSS MULTIPLY.
683X=11*315
683X=3465
X=3465/683
X=5.07 HOURS.

Answer by Alan3354(69443)

tammy drove 693 miles in 11hrs at the same rate how long would it take her to dive 315 miles?
-------------
693/11 = 315/x
693x = 3465
x = 5 hours
================
Or
693/11 = 63 mi/hr
315/63 = 5 hours